OpenStudy (anonymous):

Evaluate -4x^2-2y^3 when x=1 and y=-5

OpenStudy (anonymous):

Plug in x = 1 and y = -5 What is \[-4(1^{2})?\] and what is \[2((-5)^{3})?\] so what is your answer?

OpenStudy (anonymous):

the first is -4

OpenStudy (anonymous):

he other one noidea

OpenStudy (anonymous):

Keep in mind that any number\[x ^{3} = x * x * x\] so \[(-5)^{3} = -5*-5*-5 = ?\]

OpenStudy (anonymous):

-125

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Good, so then what's \[2*((-5)^{3}) = 2*-125\]

OpenStudy (anonymous):

-250

OpenStudy (anonymous):

Good, so then what's \[4(1^{2}) - 2((-5)^{3})?\]

OpenStudy (anonymous):

no idea hat looks too complicated

OpenStudy (anonymous):

you have both terms. you said for yourself that the first one is 4, and we worked out that the second one is -250, dont get intimidated by how it looks? you are really just calculating 4-(-250)